# Learning a Neural Solver for Parametric PDEs to Enhance Physics-Informed Methods

### v3

@@ B.1 Theoretical Analysis of PINNs
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Since $A$ is symmetric positive definite, it has eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$ with $\lambda_{i}>0$. To ensure convergence, we require $0<\eta<\frac{2}{\lambda_{\max}(A)}$. Choosing $\eta=\frac{c}{\lambda_{\max}(A)}$ with $20<c<$ 2, we have:
@@
where $\kappa(A)=\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$ is the condition number of $A$. By choosing $10<c<$ 1, we ensure $|1-c|<1$, and since $\kappa(A)\geq 1$, we have $11-\frac{c}{\kappa(A)}<$ 1. Thus, the convergence factor is:
@@ C.1 Helmholtz
We generate a dataset following the $1d$ static Helmholtz equation eq. 21. For $x\in[0,1[$,
@@
The solution can be analytically derived: $u(x)=\alpha\cos(\omega x+\beta)$, with $\beta=\arctan(\frac{-v_{0}}{\omega u_{0}})$, $\alpha=\frac{u_{0}}{\cos(\beta)}$ and directly computed from the PDE data. We generate $1,024$ trajectories for training and $256$ for testing with $u_{0},v_{0}\sim\mathcal{N}(0,1)$, and $\omega\sim\mathcal{U}(0.5,50)$ and compute the solution on $[0,1]$ with a spatial resolution of $256$. For training, we keep $800$ samples and use the complete dataset for the additional experiments presented in figs. 12(a) and 12(b). Moreover, we sub-sample the spatial resolution by $4$ and keep $64$ points for training.
@@ C.2 Poisson
We generate a dataset following the $1d$ static Poisson equation eq. 22 with forcing term. For $x\in[0,1[$,
@@
We chose $f$ to be a non-linear forcing term: $f(x)=\frac{\pi}{K}\sum_{i=1}^{K}a_{i}i^{2r}\sin(\pi x)$, with $a_{i}\sim\mathcal{U}(-100,100)$, we used $K=16$, $r=-0.5$, and solve the equation using a backward finite difference scheme. We generate $1,000$ trajectories with $u_{0},v_{0}\sim\mathcal{N}(0,1)$ and compute the solution on $[0,1]$ with a spatial resolution of $64$.
@@
We use a non-linear reaction-diffusion used in (Krishnapriyan et al., 2021; Toloubidokhti et al., 2024). This PDE has been shown to be a failure case for PINNs (Krishnapriyan et al., 2021). The PDE states as follows:
$\frac{\partial u(t,x)}{\partial t}-\nu\frac{\partial^{2}u(t,x)}{\partial x^{2}}-\rho u(t,x)(1-u(t,x)$ $=0,$ (23)
@@
We generate $800$ trajectories by varying $\nu$ in $[1,5]$ and $\rho$ in $[-5,5]$. Spatial resolution is $256$ and temporal resolution is $100$, which we sub-sample by $4$ for training, leading to a spatial resolution of $64\times 25$. The PDE is solved on $[0,1]^{2}$ as in (Toloubidokhti et al., 2024).
@@ C.3 Darcy Flow
The $2d$ Darcy Flow dataset is taken from (Li et al., 2020) and is commonly used in the operator learning literature (Li et al., 2023b; Goswami et al., 2022).
.$-\nabla.(a(x)\nabla u(x))$ $=f(x)\hskip 8.53581ptx\in(0,1)^{2},$ (26)
@@
For this dataset, the forcing term $f$ is kept constant $f=1$, and $a(x)$ is a piece-wise constant diffusion coefficient taken from (Li et al., 2020). We kept $1,000$ trajectories (on the $5,000$ available) with a spatial resolution of $64\times 64$.
@@ C.4 Heat
As proof that our method can handle $2$d + time, we consider the dataset proposed by (Zhou & Farimani, 2024).
@@
Where $L=2$, $\nu$ is randomly chosen between $[2\times 10^{-3},2\times 10^{-2}]$, $A_{j}$ in $[-0.5,0.5]$, $l_{xj},l_{xy}$ are integers in $\{1,2,3\}$ and $\phi$ is in $[0,2\pi]$. As a difference with (Zhou & Farimani, 2024), we randomly chose $J$ between $1$ and $J_{max}=5$ to have more diversity in the represented frequencies in the data. The PDEs are sampled with spatial resolution $64$ in x and y and temporal resolution $100$. However, during training the spatial resolution are subsampled by $4$ and the coordinates are re scaled between 0 and 1. As for other PDEs, we use $800$ trajectories for training and $200$ for testing.
@@ C.5 Additional dataset: Advection
The dataset is taken from (Takamoto et al., 2023).
@@
Where $\beta$ is a constant advection speed, and the initial condition is $u_{0}(x)=\sum_{k_{i}=k_{1}...k_{N}}A_{i}\sin(k_{i}x+\phi_{i})$, with $k_{i}=\frac{2\pi{n_{i}}}{L_{x}}$ and ${n_{i}}$ are randomly selected in $[1,8]$. The author used $N=2$ for this PDE. Moreover, $A_{i}$ and $\phi_{i}$ are randomly selected in $[0,1]$ and $(0,2\pi)$ respectively. Finally, $L_{x}$ is the size of the domain (Takamoto et al., 2023).
@@ C.6 Summary of problem settings considered
A summary of the datasets and parameters changing between $2$ trajectories is presented in table 3.
@@
Heat $\nu$ $[2\times 10^{-3},2\times 10^{-2}]$
@@ Appendix D Implementation details
We implemented all experiments with PyTorch (Ansel et al., 2024). We estimate the computation time needed for development and the different experiments to be approximately $300$ days.
@@ D.1 B-Spline basis
We chose to use a B-Spline basis to construct the solution. We manually build the spline and compute its derivatives thanks to the formulation and algorithms proposed in (Piegl & Tiller, 1996). We used Splines of degree $d=3$ and constructed the Splines with $2$ different configurations:
  • Take $N+d+1$ equispaced nodes of multiplicity $1$ from $\frac{d}{N}$ to $1+\frac{d}{N}$. This gives a smooth local basis with no discontinuities (see fig. 10(a)) represented by a shifted spline along the x-axis (denoted as shifted in the following).
  • Use $N+1-d$ nodes of multiplicity $1$ and $2$ nodes of multiplicity $d$ (typically on the boundary nodes: 00 and $1$). This means that nodes 00 and $1$ are not differentiable (see fig. 10(b)). We call this set-up equispaced.
@@
For $1$d + time, $2$d dataset, and $2$d + time we build a $2$ (or $3$) dimensional B-spline basis, i.e., we treat the time coordinates as spatial ones. To build such bases, we compute the Cartesian product between the $2$ (or $3$) $1$d-bases, $1$ per dimension. This means that for a $2$d dataset, for which we chose to use bases of size $N1$ and $N2$ for each coordinate, the resulting basis will have $N1+N2+N1*N2$ terms. This method makes the training more costly and several techniques to improve its scalability could be used. For $3$d datasets, the number of terms in the basis is cubic.
@@ D.2 Training details
In our experiments, neural networks are trained using the Adam optimizer. For network optimization, we employ a smooth $l1$-loss for our solver while for other baselines, we use MSE loss and/or physical losses. All models are trained for at least $1,500$ epochs on datasets composed of some sampling of $\gamma$ and/or $f$ and/or $g$. If not stated otherwise, we train our proposed method for $750$ epochs and baselines for $1,500$ epochs. We use the Adam optimizer and an initial learning rate of $0.001$. We use an exponential learning rate scheduler that lowers the learning rate by $0.995$ every epoch. Experiments are conducted on NVIDIA TITAN V ($12$ Go) for $1$d datasets and NVIDIA RTX A6000 GPU with 49Go for $1d$ + time, $2$d datasets or $2$d + time. We recall in algorithm 3 the pseudo-code for training our proposed method.
@@ D.3 Models
We present here the training details for our model and the baseline. The results can be found in table 2. In all our experiments, we use a GeLU (Hendrycks & Gimpel, 2016) activation function. The details of the model architecture on each dataset are presented in table 4.
@@
The conditioning of the problem highly depends on the parameters of the PDE and initial/boundary conditions. This can lead to unstable training when optimizing PINNs, requiring specific parameter configurations for each PDE. To avoid extensive research of the best training strategy, we found a configuration that allowed a good fitting of most of the testing dataset. This means that in the values reported in table 2, we removed trainings for which the losses exploded (only a few hard PDEs were removed, typically between none to $20$). This prevents us from extensive hyper-parameter tuning on each PDE. Please note that this lowers the reported relative MSE, thus advantaging the baseline. We detail in tables 5 and 6 the hyper-parameters for each dataset.
@@ E.1 Ablation
On fig. 11, we show that having more optimizer steps allows for a better generalization. However, we observed in our experiments that the generalization error stabilizes or even increases after the proposed $5$ steps. This limitation of the solver should be investigated in future work in order to allow the model to make more iterations. Note that for this experiment, we lowered batch size (and adapted the learning rate accordingly) when the number of steps increased (fig. 11).
@@
On figs. 12(a) and 12(b), we show that compared to other physics-informed baselines, our solver requires less data to learn to solve PDE. Note that in this simple example, the MLP+basis baseline also performed well. However, as shown in table 2, this is not the case for all the datasets. The contribution of the iterative procedure clearly appears since with 2×2$\times$2 × less data, our model performs better than this baseline.
@@
Table 7 shows that optimizing our network with physical loss greatly complicates training. Indeed, adding an ill-conditioned loss to the standard MSE makes training ill-conditioned.
@@
In table 8, we compare two optimizer configurations. "Direct" means that the network directly predicts the parameters for the next step (i.e. $\Theta_{k+1}=\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\Theta_{k}),\gamma,f,g)$), while "GD", corresponds to the update rule described before and using SGD as the base algorithm $\Theta_{k+1}=\Theta_{k}-\eta\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\Theta_{k}),\;\gamma,f,g)$. The latter clearly outperforms the direct approach and shows that learning increments is more efficient than learning a direct mapping between two updates. As shown in table 8 increasing the number of steps improves the performance (shown here for 1 and 5 update steps). However, the performance does not improve anymore after a few steps (not shown here).
@@
In table 9, we study the performance of our proposed method with different inner learning rates $\eta$. As expected, a higher learning rate leads to better performances since the optimization is taking bigger steps.
@@
As indicated in eq. 9 the inputs of our learned solver are $(\gamma,f,g,\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{L}))$. We performed experiments by removing either $\gamma$ or $\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{l})$ from the input (For the Helmholtz equation, there is no forcing term $f$). The BC $g$ are kept since they are part of the PDE specification and are required to ensure the uniqueness of the solution. This experiment (table 10) illustrates that conditioning on the PDE parameters $\gamma$ is indeed required to solve the parametric setting. Without $\gamma$, the solver has no hint on which instance should be used. The addition of the gradient information, $\nabla$, which is at the core of our method, is also crucial for improving the convergence and validates our setting.
@@
We showed by a simple experiment that our model can handle nonlinear cases (see table 11). We propose to model the solution $u$ using a non-linear combination of the basis terms $\phi_{i}$. The relation between the $\phi_{i}$ is modeled using a simple NN with one hidden layer and a tanh activation function. This experiment is performed on the Poisson PDE.
@@
Network architecture We show in table 12, an ablation on the layer type used for $\mathcal{F}_{\varrho}$ in our experiments: MLP, Residual Network (ResNet), FNO and a modified version of a MLP (ModMLP) taken from (Wang et al., 2021a), inspired from attention mechanism. We conducted this experiment on the Helmholtz dataset.
@@
Irregular grids We show in table 13, an ablation on different types of grids: regular as in table 2, and irregular. The latter were created by sampling uniformly 75% of the points in the original grid. The sampled grids are different between each trajectory both during the training and testing phases. We conducted this experiment on the Helmholtz dataset. Finally, we show some reconstruction examples in figs. 13(a) and 13(b).
@@
This experiment shows that our method is capable of handling irregular grids. We observe a decrease in the performances (table 13) and in the reconstruction quality figs. 13(a) and 13(b). However, we also note that our method is still capable of reconstructing the dynamic of the PDE, where most of the considered baselines failed to capture the oscillation in the Helmholtz PDE solution.
Error as a function of the PDE parameters values We illustrate in fig. 14 the behavior of the reconstruction of the MSE varying the PDE parameter values. We conducted this experiment on the Helmholtz PDE and varied $\omega$ from $-5$ to $55$ i.e. extrapolation of $10\%$ beyond the parameter distribution (and kept fixed boundary conditions).
@@
We observe in fig. 14 that, as expected, solving Helmholtz PDE outside the training distribution fails for our model as well as for all baselines. The proposed method performs well inside the training distribution (as already observed in table 2) and behaves similarly to fully supervised methods on Out-Of-Distribution (OOD) examples. Finally, other baselines, involving physical only or hybrid training, only predict a mean solution.

### v5

@@ B.1 Theoretical Analysis of PINNs
$\mathcal{L}_{\textnormal{Res}}(\Theta)=\frac{1}{2}\int_{\Omega}\left(\mathcal{D}u_{\Theta}(x)-f(x)\right)^{2}dx,\quad\mathcal{L}_{\textnormal{BC}}(\Theta)=\frac{1}{2}\int_{\partial\Omega}\left(u_{\Theta}(x)-g(x)\right)^{2}dx,$
@@
$A_{i,j}=\int_{\Omega}\left(\mathcal{D}\phi_{i}(x)\right)\left(\mathcal{D}\phi_{j}(x)\right)dx+\lambda\int_{\partial\Omega}\phi_{i}(x)\phi_{j}(x)dx.$ (15)
@@
$\nabla_{\Theta}\mathcal{L}_{\textnormal{Res}}(\Theta)=\int_{\Omega}\left(\mathcal{D}u_{\Theta}(x)-f(x)\right)\mathcal{D}\phi(x)\,dx,$
@@
$\nabla_{\Theta}\mathcal{L}_{\textnormal{BC}}(\Theta)=\int_{\partial\Omega}\left(u_{\Theta}(x)-g(x)\right)\phi(x)\,dx,$
@@
$A_{i,j}$ $=\int_{\Omega}\left(\mathcal{D}\phi_{i}(x)\right)\left(\mathcal{D}\phi_{j}(x)\right)dx+\lambda\int_{\partial\Omega}\phi_{i}(x)\phi_{j}(x)dx,$ (16)
$b_{i}$ $=\int_{\Omega}f(x)\mathcal{D}\phi_{i}(x)\,dx+\lambda\int_{\partial\Omega}g(x)\phi_{i}(x)\,dx.$
@@
Since $A$ is symmetric positive definite, it has eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$ with $\lambda_{i}>0$. To ensure convergence, we require $0<\eta<\frac{2}{\lambda_{\max}(A)}$. Choosing $\eta=\frac{c}{\lambda_{\max}(A)}$ with $0<c<2$, we have:
@@
where $\kappa(A)=\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$ is the condition number of $A$. By choosing $0<c<1$, we ensure |1c|<1|1-c|<1, and since $\kappa(A)\geq 1$, we have $1-\frac{c}{\kappa(A)}<1$. Thus, the convergence factor is:
@@ C.1 Helmholtz
We generate a dataset following the $1d$ static Helmholtz equation eq.˜21. For $x\in[0,1[$,
@@
The solution can be analytically derived: $u(x)=\alpha\cos(\omega x+\beta)$, with $\beta=\arctan(\frac{-v_{0}}{\omega u_{0}})$, $\alpha=\frac{u_{0}}{\cos(\beta)}$ and directly computed from the PDE data. We generate $1,024$ trajectories for training and $256$ for testing with $u_{0},v_{0}\sim\mathcal{N}(0,1)$, and $\omega\sim\mathcal{U}(0.5,50)$ and compute the solution on $[0,1]$ with a spatial resolution of $256$. For training, we keep $800$ samples and use the complete dataset for the additional experiments presented in figs.˜12(a) and 12(b). Moreover, we sub-sample the spatial resolution by $4$ and keep $64$ points for training.
@@ C.2 Poisson
We generate a dataset following the $1d$ static Poisson equation eq.˜22 with forcing term. For $x\in[0,1[$,
@@
We chose $f$ to be a non-linear forcing term: $f(x)=\frac{\pi}{K}\sum_{i=1}^{K}a_{i}i^{2r}\sin(\pi x)$, with $a_{i}\sim\mathcal{U}(-100,100)$, we used $K=16$, r=0.5r=-0.5, and solve the equation using a backward finite difference scheme. We generate $1,000$ trajectories with $u_{0},v_{0}\sim\mathcal{N}(0,1)$ and compute the solution on $[0,1]$ with a spatial resolution of $64$.
@@
We use a non-linear reaction-diffusion used in (krishnapriyan_characterizing; toloubidokhti2024dats). This PDE has been shown to be a failure case for PINNs (krishnapriyan_characterizing). The PDE states as follows:
$\frac{\partial u(t,x)}{\partial t}-\nu\frac{\partial^{2}u(t,x)}{\partial x^{2}}-\rho u(t,x)(1-u(t,x))$ $=0,$ (23)
@@
We generate $800$ trajectories by varying $\nu$ in $[1,5]$ and $\rho$ in [5,5][-5,5]. Spatial resolution is $256$ and temporal resolution is $100$, which we sub-sample by $4$ for training, leading to a spatial resolution of $64\times 25$. The PDE is solved on $0,1]^{2}$ as in (toloubidokhti2024dats).
@@ C.3 Darcy Flow
The $2d$ Darcy Flow dataset is taken from (li2020fno) and is commonly used in the operator learning literature (li2023pino; goswami2022PIDON).
$-\nabla.(a(x)\nabla u(x))$ $=f(x)\hskip 8.53581ptx\in(0,1)^{2},$ (26)
@@
For this dataset, the forcing term $f$ is kept constant $f=1$, and $a(x)$ is a piece-wise constant diffusion coefficient taken from (li2020fno). We kept $1,000$ trajectories (on the $5,000$ available) with a spatial resolution of $64\times 64$.
@@ C.4 Heat
As proof that our method can handle 22d + time, we consider the dataset proposed by (maepde24).
@@
Where $L=2$, $\nu$ is randomly chosen between $2\times 10^{-3},2\times 10^{-2}]$, $A_{j}$ in [0.5,0.5][-0.5,0.5], $l_{xj},l_{xy}$ are integers in $\{1,2,3\}$ and $\phi$ is in $0,2\pi]$. As a difference with (maepde24), we randomly chose $J$ between $1$ and $J_{max}=5$ to have more diversity in the represented frequencies in the data. The PDEs are sampled with spatial resolution $64$ in x and y and temporal resolution $100$. However, during training the spatial resolution are subsampled by $4$ and the coordinates are re scaled between 0 and 1. As for other PDEs, we use $800$ trajectories for training and $200$ for testing.
@@ C.5 Additional dataset: Advection
The dataset is taken from (takamoto2023pdebench).
@@
Where $\beta$ is a constant advection speed, and the initial condition is $u_{0}(x)=\sum_{k_{i}=k_{1}...k_{N}}A_{i}\sin(k_{i}x+\phi_{i})$, with $k_{i}=\frac{2\pi{n_{i}}}{L_{x}}$ and $n_{i}$} are randomly selected in $[1,8]$. The author used $N=2$ for this PDE. Moreover, $A_{i}$ and $\phi_{i}$ are randomly selected in $[0,1]$ and $(0,2\pi)$ respectively. Finally, $L_{x}$ is the size of the domain (takamoto2023pdebench).
@@ C.6 Summary of problem settings considered
A summary of the datasets and parameters changing between $2$ trajectories is presented in table˜3.
@@
Heat $\nu$ $2\times 10^{-3},2\times 10^{-2}]$
@@ Appendix D Implementation details
We implemented all experiments with PyTorch (PyTorch2). We estimate the computation time needed for development and the different experiments to be approximately $300$ days.
@@ D.1 B-Spline basis
We chose to use a B-Spline basis to construct the solution. We manually build the spline and compute its derivatives thanks to the formulation and algorithms proposed in (PiegTill96_NURBS). We used Splines of degree $d=3$ and constructed the Splines with $2$ different configurations:
  • Take $N+d+1$ equispaced nodes of multiplicity $1$ from $\frac{d}{N}$ to $1+\frac{d}{N}$. This gives a smooth local basis with no discontinuities (see fig.˜10(a)) represented by a shifted spline along the x-axis (denoted as shifted in the following).
  • Use N+1dN+1-d nodes of multiplicity $1$ and $2$ nodes of multiplicity $d$ (typically on the boundary nodes: 0 and $1$). This means that nodes 0 and $1$ are not differentiable (see fig.˜10(b)). We call this set-up equispaced.
@@
For 11d + time, 22d dataset, and 22d + time we build a $2$ (or $3$) dimensional B-spline basis, i.e., we treat the time coordinates as spatial ones. To build such bases, we compute the Cartesian product between the $2$ (or $3$) 11d-bases, 11 per dimension. This means that for a 22d dataset, for which we chose to use bases of size $N1$ and $N2$ for each coordinate, the resulting basis will have N1+N2+N1N2N1+N2+N1*N2 terms. This method makes the training more costly and several techniques to improve its scalability could be used. For 33d datasets, the number of terms in the basis is cubic.
@@ D.2 Training details
In our experiments, neural networks are trained using the Adam optimizer. For network optimization, we employ a smooth $l1$-loss for our solver while for other baselines, we use MSE loss and/or physical losses. All models are trained for at least $1,500$ epochs on datasets composed of some sampling of $\gamma$ and/or $f$ and/or $g$. If not stated otherwise, we train our proposed method for $750$ epochs and baselines for $1,500$ epochs. We use the Adam optimizer and an initial learning rate of $0.001$. We use an exponential learning rate scheduler that lowers the learning rate by $0.995$ every epoch. Experiments are conducted on NVIDIA TITAN V ($12$ Go) for 11d datasets and NVIDIA RTX A6000 GPU with 49Go for $1d$ + time, 22d datasets or 22d + time. We recall in algorithm˜3 the pseudo-code for training our proposed method.
@@ D.3 Models
We present here the training details for our model and the baseline. The results can be found in table˜2. In all our experiments, we use a GeLU (Hendrycks2016_GELU) activation function. The details of the model architecture on each dataset are presented in table˜4.
@@
The conditioning of the problem highly depends on the parameters of the PDE and initial/boundary conditions. This can lead to unstable training when optimizing PINNs, requiring specific parameter configurations for each PDE. To avoid extensive research of the best training strategy, we found a configuration that allowed a good fitting of most of the testing dataset. This means that in the values reported in table˜2, we removed trainings for which the losses exploded (only a few hard PDEs were removed, typically between none to $20$). This prevents us from extensive hyper-parameter tuning on each PDE. Please note that this lowers the reported relative MSE, thus advantaging the baseline. We detail in tables˜5 and 6 the hyper-parameters for each dataset.
@@ E.1 Ablation
On fig.˜11, we show that having more optimizer steps allows for a better generalization. However, we observed in our experiments that the generalization error stabilizes or even increases after the proposed $5$ steps. This limitation of the solver should be investigated in future work in order to allow the model to make more iterations. Note that for this experiment, we lowered batch size (and adapted the learning rate accordingly) when the number of steps increased (fig.˜11).
@@
On figs.˜12(a) and 12(b), we show that compared to other physics-informed baselines, our solver requires less data to learn to solve PDE. Note that in this simple example, the MLP+basis baseline also performed well. However, as shown in table˜2, this is not the case for all the datasets. The contribution of the iterative procedure clearly appears since with $2\times$ less data, our model performs better than this baseline.
@@
Table˜7 shows that optimizing our network with physical loss greatly complicates training. Indeed, adding an ill-conditioned loss to the standard MSE makes training ill-conditioned.
@@
In table˜8, we compare two optimizer configurations. "Direct" means that the network directly predicts the parameters for the next step (i.e. $\Theta_{k+1}=\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\Theta_{k}),\gamma,f,g)$), while "GD", corresponds to the update rule described before and using SGD as the base algorithm $\Theta_{k+1}=\Theta_{k}-\eta\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\Theta_{k}),\;\gamma,f,g)$. The latter clearly outperforms the direct approach and shows that learning increments is more efficient than learning a direct mapping between two updates. As shown in table˜8 increasing the number of steps improves the performance (shown here for 1 and 5 update steps). However, the performance does not improve anymore after a few steps (not shown here).
@@
In table˜9, we study the performance of our proposed method with different inner learning rates $\eta$. As expected, a higher learning rate leads to better performances since the optimization is taking bigger steps.
@@
As indicated in eq.˜9 the inputs of our learned solver are $(\gamma,f,g,\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{L}))$. We performed experiments by removing either $\gamma$ or $\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{l})$ from the input (For the Helmholtz equation, there is no forcing term $f$). The BC $g$ are kept since they are part of the PDE specification and are required to ensure the uniqueness of the solution. This experiment (table˜10) illustrates that conditioning on the PDE parameters $\gamma$ is indeed required to solve the parametric setting. Without $\gamma$, the solver has no hint on which instance should be used. The addition of the gradient information, $\nabla$, which is at the core of our method, is also crucial for improving the convergence and validates our setting.
@@
We showed by a simple experiment that our model can handle nonlinear cases (see table˜11). We propose to model the solution $u$ using a non-linear combination of the basis terms $\phi_{i}$. The relation between the $\phi_{i}$ is modeled using a simple NN with one hidden layer and a tanh activation function. This experiment is performed on the Poisson PDE.
@@
Network architecture We show in table˜12, an ablation on the layer type used for $\mathcal{F}_{\varrho}$ in our experiments: MLP, Residual Network (ResNet), FNO and a modified version of a MLP (ModMLP) taken from (wang2021understanding), inspired from attention mechanism. We conducted this experiment on the Helmholtz dataset.
@@
Irregular grids We show in table˜13, an ablation on different types of grids: regular as in table˜2, and irregular. The latter were created by sampling uniformly 75% of the points in the original grid. The sampled grids are different between each trajectory both during the training and testing phases. We conducted this experiment on the Helmholtz dataset. Finally, we show some reconstruction examples in figs.˜13(a) and 13(b).
@@
This experiment shows that our method is capable of handling irregular grids. We observe a decrease in the performances (table˜13) and in the reconstruction quality figs.˜13(a) and 13(b). However, we also note that our method is still capable of reconstructing the dynamic of the PDE, where most of the considered baselines failed to capture the oscillation in the Helmholtz PDE solution.
Error as a function of the PDE parameters values We illustrate in fig.˜14 the behavior of the reconstruction of the MSE varying the PDE parameter values. We conducted this experiment on the Helmholtz PDE and varied $\omega$ from 5-5 to $55$ i.e. extrapolation of $10\%$ beyond the parameter distribution (and kept fixed boundary conditions).
@@
We observe in fig.˜14 that, as expected, solving Helmholtz PDE outside the training distribution fails for our model as well as for all baselines. The proposed method performs well inside the training distribution (as already observed in table˜2) and behaves similarly to fully supervised methods on Out-Of-Distribution (OOD) examples. Finally, other baselines, involving physical only or hybrid training, only predict a mean solution.